OpenStudy (anonymous):

Identify the x and y intercepts for the equation 4x + 3y = 12

OpenStudy (anonymous):

To solve for the x-intercept, set y=0 To solve for the y-intercept, set x=0

OpenStudy (anonymous):

so now what ?

OpenStudy (anonymous):

4x+3(0)=12 4(0)+3y=12 Simplify and the answers and will be your x and y-intercepts

OpenStudy (tkhunny):

Even cooler, just divide by 12. \(\dfrac{x}{3} + \dfrac{y}{4} = 1\) And the intercepts are just staring at you.

OpenStudy (anonymous):

im so lost

OpenStudy (anonymous):

?

OpenStudy (tkhunny):

That form I showed you is called the "Intercept Form". Just read the denominators. x intercept (3,0) y intercept (0,4) That's it.